OpenStudy (anonymous):

solve system 4a+2b-24=0 10a-2b-56=0 DUE IN 10 MIN!!!!!!!

OpenStudy (anonymous):

Use elimination for b Solve for a using the value you get for b done

OpenStudy (anonymous):

i've tried multiple times with all wrong and different answers

OpenStudy (anonymous):

What did you get for a value when you eliminated b?

OpenStudy (anonymous):

a=40/7

OpenStudy (anonymous):

i got 40/7=a and 4/7=b it says its wrong

OpenStudy (wolfe8):

Careful with signs

OpenStudy (anonymous):

Yeah wolfe is right. You're not being aware of your subtraction with negative integers.

OpenStudy (anonymous):

-24 - 56 = ?

OpenStudy (anonymous):

-80

OpenStudy (anonymous):

divide that by 14 a = -80/14 simplify that

OpenStudy (anonymous):

-40/7

OpenStudy (anonymous):

i tried this also

OpenStudy (anonymous):

you can't simplify it even more so...

OpenStudy (anonymous):

what about for b

OpenStudy (anonymous):

i have 3 min. can i pls get an answer just this once

OpenStudy (anonymous):

well when you substitute a into let's say equation 1: you get 164/7

OpenStudy (anonymous):

that's all i know

OpenStudy (anonymous):

what about b

OpenStudy (anonymous):

that's what i just told you

OpenStudy (anonymous):

its ok time ran out

OpenStudy (anonymous):

oh..sorry